# Testing the Equivalence Principle on Cosmological Scales Using Peculiar Acceleration Power Spectrum

### v1

@@ Appendix A Validation of acceleration power spectrum
Figure 3: Comparison of halo acceleration power spectra from simulations and linear theory predictions for four redshift snapshots: low-mass halos (left) and high-mass halos (right).
@@
$-k^{2}\phi(\bm{k},z)=\frac{3}{2}\Omega_{m}H^{2}a^{2}\delta_{m}\,.$ (12)
@@
$=$ $\frac{3}{4}\left(\frac{H^{2}_{0}\Omega_{m0}}{a^{2}k^{2}}\right)^{2}P_{m}(k,z)\,.$
@@
We also observe that, on scales $k\lesssim 0.3~{}h/\rm Mpc$, the power spectra for the two halo mass bins at the same redshift snapshot are nearly identical. This indicates that $E_{\rm ep}$ remains robust across this scale range. A more detailed discussion of this result is provided in Sect. III.

### v2

@@ Appendix A Validation of acceleration power spectrum
Refer to captionRefer to caption
Figure 3: Comparison of halo acceleration power spectra from simulations and linear theory predictions for four redshift snapshots: small-mass halos (left) and large-mass halos (right).
@@
$k^{2}\phi(\bm{k},z)=\frac{3}{2}\Omega_{m}H^{2}a^{2}\delta_{m}\,.$ (14)
@@
$=$ $\frac{3}{4}\left(\frac{H^{2}_{0}\Omega_{m0}}{a^{2}k}\right)^{2}P_{m}(k,z)\,.$
@@
We also observe that, on scales $k\lesssim 0.2~h/\rm Mpc$, the power spectra for the two halo mass bins at the same redshift snapshot are nearly identical. This indicates that $E_{\rm ep}$ remains robust across this scale range. A more detailed discussion of this result is provided in Sect. III.